A ball is projected from a point on the floor with a speed of 15 m/s at an angel of 60° with the horizontal. Will it hit a vertical wall 5 m away from the point of projection and perpendicular to the plane of projection without hitting the floor?
Will the answer differ if the wall is 22 m away?

Grade:11

1 Answers

Kevin Nash
9 years ago
Sol. Here u = 15 m/s, θ = 60°, g = 9.8 m/s2 Horizontal range X = (u^2 sin⁡2θ)/g = ((15)^2 sin⁡〖(2 x 60°)〗)/9.8 = 19.88 m In first case the wall is 5 m away from projection point, so it is in the horizontal range of projectile. So the ball will hit the wall. In second case (22 m away) wall is not within the horizontal range. So the ball would not hit the wall.